OpenStudy (anonymous):

13+5 over x=6 is x3?

OpenStudy (anonymous):

what does the problem say?

jimthompson5910 (jim_thompson5910):

If you mean to write \[\Large \frac{15+3}{x} = 6\] then x = 3 is correct.

OpenStudy (anonymous):

13+5 over 3 (theres a line between the top and bottom =6 .. is x 3?

OpenStudy (anonymous):

\[\large 13+\frac{5}{x}=6\]?

jimthompson5910 (jim_thompson5910):

oops

OpenStudy (anonymous):

no like this

jimthompson5910 (jim_thompson5910):

If you mean to write \[\Large \frac{13+5}{x} = 6\] then x = 3 is correct.

OpenStudy (anonymous):